# Reliability in Semantic Segmentation: Can We Use Synthetic Data?

### v1

@@ 3.3 Confidence Calibration with Synthetic Data
Synthetic Images Semantic-FPN MobileNetV3 SegF-B5
@@
We perform temperature scaling using our synthetic data. Temperature scaling [7] is a well-established technique for calibrating pretrained models, typically conducted on a small validation set within the target domain. In our study, for each segmenter, we utilized the same sets of synthetic data generated in Section 3.2 to optimize temperature scaling factors, with one adjustment made for each covariate shift. For comparison, we replicate the process using real-shift data from ACDC and IDD.
Figure 6 reports the calibration improvement for the $40$ pretrained segmenters using either real-shift data ( ) or synthetic data ( ). The subplots are arranged in increasing domain gap order from top to bottom, with the segmenters ranked from left to right based on increasing robustness on real-shift data. The Expected Calibration Error (ECE) quantifies the calibration results, with a lower ECE indicating a better-calibrated model. For better visualization and interpretation, we present the relative ECE improvement, computed as the percentage decrease in ECE after calibration compared to the original ECE without calibration. For example, a model with an ECE of 0.4 before calibration and an ECE of 0.2 after calibration will achieve a $(0.4-0.2)/0.4=50\%$ relative improvement.
We observe promising calibration results when employing our synthetic data. While not as good as real-shift data, synthetic data achieves a promising success rate of $72.5\%$ on IDD and exceeds $90\%$ on the four ACDC shifts. Interestingly, in weather shifts, we empirically observe that more robust models derive greater benefits when calibrated using our synthetic data; the reverse is observed for ‘europe-india’ geographical shift. While with real-shift data, robustness and calibration are not well correlated [5], our results suggest that a potential correlation might exist between the two factors when using synthetic data. In Appendix B, we provide technical details on our calibration approaches and additional results.
@@ 3.4 On Practical Applicability
One significant advantage of our framework lies in its potential to address rare conditions simply through prompting. The practical applicability of generative testing is tremendous. Our results demonstrate promising signals; practitioners can begin assessing and ranking their pretrained models for a new, unseen target domain of interest without the need for real data collection. In practice, our proposed generative benchmarking can serve as the initial step in a full validation pipeline, helping filter out non-robust prototypes and thereby saving on total operational costs. Starting from complementing real-data validation, one can envision a future where generative techniques mature to the point of fully replacing real-data validation. In Figure 7, we visualize some synthetic images and model predictions under rare conditions, such as being flooded with water, having autumn leaves scattered across the road, or having a building on fire. We observe clear visual distinctions between weaker (Semantic-FPN and MobileNetV3) and stronger (SegF-B5) models, knowing that their Cityscapes scores do not differ significantly. More examples are provided in Appendix C.
@@ 4.1 Inpainting Anomaly Objects
We inpaint random objects into Cityscapes images. To this end, we initially sample a location —- a square box to which we inpaint the new object. We crop the box, upsample its content to match the preferred output size of the generative model, and inpaint an object guided by a text prompt. In this step, we leverage Stable Diffusion inpainting capabilities, obtaining high-definition square images of the desired object. This image is then resized and pasted back into the original image, creating a final high-definition synthetic image. To ensure compositional consistency, we employ two techniques: Firstly, we divide the cropped box into two regions by center cropping it again. We inpaint only the inner region, leaving the outer region untouched, similar to the approach in RePaint [19]. Secondly, after composing the final image, we address any remaining inconsistencies by applying a light noise over the entire picture and performing reverse diffusion again. Technical details and visualizations are provided in Appendix A.
@@
Figure 8: Entropy Correlation. The top-left inset reports the Pearson correlations between real-OOD entropy vs. synthetic-OOD entropy, computed either on ‘curated‘ ( ) or all synthetic inpainted images ( ). Evaluations are performed on the same model set used in Figure 3, with similar markers.
[…]
@@
Figure 10: Qualitative results. Confidence maps are visualized for the four exemplified models on real data (first row) and synthetic inpainted data (second and third rows). Hotter colors correspond to higher OOD likelihood. Ideally, results should exhibit hot colors in OOD areas and cold colors everywhere else. We observe a strong correlation in model reactions to real and synthetic OOD regions, particularly for more recent and robust models.
Our end-to-end generation pipeline is fully automatic. Through qualitative assessment, we achieve a satisfactory success rate in terms of generation realism; some inpainted images are illustrated in Figure 10 and much more in Appendix C. However, this still leaves a few generations with artifacts, characterized by either unusual compositions or unrealistic details. We here question the criticality of realism in assessing OOD detection and, furthermore, in improving OOD detection. To this end, we construct two different sets: (i) all 23,040 images generated automatically and (ii) 656 curated images where we manually select the best images in terms of visual quality and realism. We note that the manual selection process for the curated set is not exhaustive and is constrained by our allocated resources; there are many more high-quality images in the ‘all’ set. In what follows, we present results using both curated and uncurated sets.
@@
4.2 Assess OOD Detection
Experimental setup. To measure how the segmenters react to unseen OOD objects, we use standard anomaly detection metrics [36], that are False Positive Rate at 95% true positives ($\text{FPR}_{\text{95}}$), Area Under ROC curve (AUROC), and Area Under Precision-Recall curve (AUPR). AUPR are declined into $\text{AUPR}_{\text{IN}}$ and $\text{AUPR}_{\text{OUT}}$, that considers the in-distribution regions, respectively the out-of-distribution regions (the inpainted object), as positive regions to compute the Precision-Recall curves.
All segmenters in our study are not designed to produce confidence scores. We thus seek for various techniques to derive confidence scores from pretrained models [12] and eventually narrow down the options to two measures: (i) Entropy of soft-probability predictions, and (ii) MaxLogit as the maximum logit value (before softmax) among the classes. While Entropy is the traditional measure of uncertainty, MaxLogit is a recent and surprising finding that has been proven to be more effective in estimating OOD confidence [12].
Results. For quantitative comparison, we leverage the SegmentMeIfYouCan (SMIYC) dataset [4], a recent dataset for OOD detection. We resort to the RoadAnomaly21 split in SMIYC, due to similarity in object scales to our synthetic data. We analyze the correlation between the OOD scores obtained on RoadAnomaly21 and one using our synthetic inpainted data.  Figure 8 reports our first analysis on the entropy in the OOD areas, either real (RoadAnomaly21) or generated. For each model, we compute the Pearson Correlation (PCC) between real-OOD entropy and synthetic-OOD entropy; the computation is done on both ‘curated‘ and ‘all‘ sets. In terms of entropy, we observe a very high correlation between real- and synthetic-OOD, reaching $0.94$ PCC using both ‘curated‘ and ‘all‘ sets.
We then analyze the correlation between real and synthetic anomaly metrics.  Figure 9 presents our primary findings and Figure 10 illustrates some qualitative results. We observe a strong correlation with real scores when utilizing the ’curated’ set for computing synthetic scores; the curated PCCs ( ) are consistently around 0.8 across multiple metrics, irrespective of the two confidence measures. Although the correlations are somewhat weaker when using all uncurated synthetic data ( ), such results remain acceptable, particularly when no effort is dedicated to curation.
@@ 4.3 Improve OOD Detection
OOD metrics are computed on RoadAnomaly21 and reported in Table 1. The RbA models trained on our data significantly outperform the vanilla RbA model. We reach comparable performance to the RbA variants that leverage the external COCO dataset for augmentation. Notably, there are no clear differences between using ‘curated’ or ‘all’ sets. We conjecture that, unlike benchmarking, training for OOD detection does not demand a high degree of realism from synthetic data. This explains why the simple strategy of copy-pasting COCO objects already proves effective. All results are consistent across the two addressed backbones.  Figure 10 illustrates a few qualitative results.
@@ 5 Takeaways
In this work, we explore the potential of synthetic data in reliability assessment for semantic segmentation networks. Our promising results encourage further collective investigations into this research problem, paving the way for synthetic system validation, especially in safety-critical applications. We summarize here our primary findings:
@@
    Reliability Under Covariate Shifts: synthetic data can help assess the relative robustness of models in real-life covariate shifts, especially useful when shifts to the training condition are significant. Synthetic data can well complement real data in system validation, helping reduce the total operational cost. Pretrained models can be calibrated using synthetic data to better estimate prediction confidence in any shifted domains.
@@
Limitations. We focus our research on the task of semantic segmentation while keeping open the extension possibility to other critical tasks, such as object detection. Our quantitative assessments are confined to existing publicly available datasets. However, our framework is fully zero-shot and can be applied to any domain of interest. On the generative side, our study is restricted to Stable Diffusion and ControlNet due to our resource constraints. Of note, although improvements in this area should enhance the results, similar insights are expected to be achieved.
Acknowledgements. This work is supported by ELSA - European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. We thank the authors of RELIS [5] for providing us their pretrained checkpoints.
Appendix
In this appendix, we provide technical details for covariate shifts generation and OOD object inpainting (Appendix A), additional calibration details and results (Appendix B), and more qualitative examples (Appendix C).
Appendix A Technical Details
A.1 Covariate Shifts Training
@@
A.2 Covariate Shifts Generation
To generate images with new styles, we take a semantic mask from the validation set of Cityscapes, crop and resize it as explained in the previous part. We use nearest neighbor interpolation to keep good values for specific classes. We only use the part of the caption extracted with CLIP-interrogator which corresponds to a BLIP caption. To this new caption, we add [, in $<$domain$>$] depending on the domain we want to generalize to. Starting from pure noise, we use 25 DDIM steps with a guidance scale of 8. On a RTX 2080, one new image is generated in about 4 seconds. All other sampling hyperparameters are the per default settings on the official ControlNet repository.
A.3 OOD Objects Generation
@@
The OOD object generation pipeline described in Section 4.1 is further illustrated in Figure 11. It is composed of three steps. Given a text prompt containing an object, the inpainting step generates a zoomed-in version of the object with the appropriate close-range background given as context. The mask extraction step infers the anomaly mask from the zoomed-in generated image and the name of the object. Both are in-pasted back in the original complete image or mask. To reduce some composition artifacts, the composite image is refined with a noise/denoise step.
In details, we first randomly choose a box size for the new object, uniformly sampled between a quarter and half the minimum dimension of the original image. We also uniformly sample a location for the box in the bottom three quarters of the image. This box will contain the new object we wish to add, and we refer to it as inpainted region. In addition to the inpainted region, we create a larger box, 1.5$\times$ its height and width, with the inpainted region in its center. The contour outside of the inpainted region will serve as context for the inpainting process. We then crop and resize the inpainted region with its context to a resolution of 512 $\times$ 512. We fully noise the inpainted region, but leave the clean context. We then denoise the inpainted zone with the prompt “A photo of an [normal-[[[object]normal-]]], with a guidance scale of 15. The full patch is then resized and pasted on the original image, at its original position. Some artifact might be still present as shown in Figure 12. To remedy this, we refine the inpainted zone by noising and denoising it with 0.65 strength, with the default guidance scale of 7.5. The effect of this refining step is shown in Figure 12.
@@ A.4 OOD Detection Training
For the OOD detection method in Section 4.3, we used the codebase of [22]. We adapt the code to be able to use our generated data with the binary masks extracted from Grounded-SAM, as explained in Section A.3. As in the original paper [22], we fine-tune the mask prediction MLP and classification layer after the transformer decoder. To obtain all results reported in Table 1, we used the recommended hyperparameters, and train the models for 5000 iterations.
A.5 Segmentation Models

### v2

@@ 3.3 Confidence Calibration with Synthetic Data
We perform temperature scaling using our synthetic data. Temperature scaling [6] is a well-established technique for calibrating pretrained models, typically conducted on a small validation set within the OOD domain. In our study, for each segmenter, we utilized the same sets of synthetic data generated in Sec. 3.2 to optimize temperature scaling factors, with one adjustment made for each covariate shift. For comparison, we replicate the process using real-shift data from ACDC and IDD.
Fig. 6 reports the calibration improvement for the $40$ pretrained segmenters using either real-shift data ( ) or synthetic data ( ). The subplots are arranged in increasing domain gap order from top to bottom, with the segmenters ranked from left to right based on increasing robustness on real-shift data. The Expected Calibration Error (ECE) [24] quantifies the calibration results, with a lower ECE indicating a better-calibrated model. For better interpretation, we present the relative ECE improvement, computed as the percentage decrease in ECE after calibration compared to the original ECE without calibration. For example, a model with an ECE of 0.4 before calibration and an ECE of 0.2 after calibration will achieve a $(0.4-0.2)/0.4=50\%$ relative improvement.
Figure 6: ECE improvement using synthetic data. The 40 segmenters are calibrated using either real-shift data or synthetic data. For each model, the relative ECE improvement (%) over its non-calibrated version is computed, visualized by   (synthetic shift) and   (real shift). The subplots are ranked by model robustness on real-shift data (left-to-right) and by the increased domain shift (top-to-bottom). The titles of the subplots indicate the percentage of models that showed improvement with synthetic data.
We observe promising calibration results when employing our synthetic data. While not as good as real-shift data, synthetic data achieves a promising success rate of $72.5\%$ on IDD and exceeds $90\%$ on the four ACDC shifts. Interestingly, in weather shifts, we empirically observe that more robust models derive greater benefits when calibrated using our synthetic data; the reverse is observed for ‘europe-india’ geographical shift. While with real-shift data, robustness and calibration are not well correlated [14], our results suggest that a potential correlation might exist between the two factors when using synthetic data. We note that calibration with temperature scaling does not always guarantee ECE improvement. Such phenomenon may happen even using real data, especially under domain shifts as explored in prior work [26]. In Appendix 0.B, we provide more technical details and results.
Syn. Img Sem.FPN MobileV3 SegF-B5
@@
Figure 7: Qualitative results. Examples of rare conditions generated for testing and predictions from different models. Results of the strong model like SegFormer-B5 are visibly better than the Semantic-FPN and MobileNetV3.
@@ 3.4 On Practical Applicability
One significant advantage of our framework lies in its potential to address rare conditions simply through prompting. The practical applicability of generative testing is tremendous. Our results demonstrate promising signals; practitioners can begin assessing and ranking their pretrained models for a new, unseen OOD domain of interest without the need for real data collection. In practice, our proposed generative benchmarking can serve as the initial step in a full validation pipeline, helping filter out non-robust prototypes and thereby saving on total operational costs. Starting from complementing real-data validation, one can envision a future where generative techniques mature to the point of fully replacing real-data validation. In Fig. 7, we visualize some synthetic images and model predictions under rare conditions, such as being flooded with water, having autumn leaves scattered across the road, or having a building on fire. We observe clear visual distinctions between weaker (Semantic-FPN and MobileNetV3) and stronger (SegF-B5) models, knowing that their Cityscapes scores do not differ significantly. More examples are provided in Appendix 0.F.
@@ 4.1 Inpainting Anomaly Objects
We inpaint random objects into Cityscapes images. To this end, we initially sample a location — a square box to which we inpaint the new object. We crop the box, upsample its content to match the preferred output size of the generative model, and inpaint an object guided by a text prompt. In this step, we leverage Stable Diffusion inpainting capabilities, obtaining high-definition square images of the desired object. This image is then resized and pasted back into the original image, creating a final high-definition synthetic image. To ensure compositional consistency, we employ two techniques: Firstly, we divide the cropped box into two regions by center cropping it again. We inpaint only the inner region, leaving the outer region untouched, similar to the approach in RePaint [21]. Secondly, after composing the final image, we address any remaining inconsistencies by applying a light noise over the entire picture and performing reverse diffusion again. Details and visualizations are provided in Appendix 0.A.
@@
Figure 8: Entropy Correlation. The top-left inset reports the Pearson correlations between real-OOD entropy vs. synthetic-OOD entropy, computed either on ‘curated‘ ( ) or all synthetic inpainted images ( ). Evaluations are performed on the same model set used in Fig. 3, with similar markers.
Figure 9: Correlation in OOD Detection. Each subplot scatters computed anomaly scores of segmenters on real OODs (y-axis) and on synthetic OODs (x-axis). The top row shows the four anomaly metrics utilized: FPR95, AUROC, AUPRIN, and AUPROUT. The results are organized into two rows corresponding to two different confidence measures (i) Entropy and (ii) MaxLogit. In the top-left corner of each subplot, an inset plots Pearson correlations to real OOD for ‘curated‘ ( ) and ‘all‘ ( ) synthetic sets. Evaluations are performed on the same model set used in Fig. 3, with similar markers.
Our end-to-end generation pipeline is fully automatic. Through qualitative assessment, we achieve a satisfactory success rate in terms of generation realism; some inpainted images are illustrated in Fig. 10 and much more in Fig. 16. However, this still leaves a few generations with artifacts, characterized by either unusual compositions or unrealistic details. We here question the criticality of realism in assessing OOD detection and, furthermore, in improving OOD detection. To this end, we construct two different sets: (i) all 23,040 images generated automatically and (ii) 656 curated images where we manually select the best images in terms of visual quality and realism; specifically, curators filter out strong color saturation differences or partial objects, e.g. animal heads. We note that the manual selection process for the curated set is not exhaustive and is constrained by our allocated resources; there are many more high-quality images in the ‘all’ set. In what follows, we present results using both curated and uncurated sets.
4.2 Assess OOD Detection
Experimental setup. To measure how the segmenters react to unseen OOD objects, we use standard anomaly detection metrics [41], which are False Positive Rate at 95% true positives ($\text{FPR}_{\text{95}}$), Area Under ROC curve (AUROC), and Area Under Precision-Recall curve (AUPR). AUPR are declined into $\text{AUPR}_{\text{IN}}$ and $\text{AUPR}_{\text{OUT}}$, which consider the in-distribution regions, respectively the out-of-distribution regions (the inpainted object), as positive regions to compute the Precision-Recall curves.
All segmenters in our study are not designed to produce confidence scores. We thus seek various techniques to derive confidence scores from pretrained models [9] and eventually narrow down the options to two measures: (i) Entropy of soft-probability predictions, and (ii) MaxLogit as the maximum logit value (before softmax) among the classes. While Entropy is the traditional measure of uncertainty, MaxLogit is a recent and surprising finding that has been proven to be more effective in estimating OOD confidence [9].
Results. For quantitative comparison, we leverage the SegmentMeIfYouCan (SMIYC) dataset [4], a recent dataset for OOD detection. We resort to the RoadAnomaly21 split in SMIYC, due to similarity in object scales to our synthetic data. We analyze the correlation between the OOD scores obtained on RoadAnomaly21 and one using our synthetic inpainted data.  Fig. 8 reports our first analysis on the entropy in the OOD areas, either real or generated. For each model, we compute the Pearson Correlation (PCC) between real-OOD entropy and synthetic-OOD entropy; the computation is done on both ‘curated‘ and ‘all‘ sets. We observe a very high entropy correlation between real- and synthetic-OOD, reaching $0.94$ PCC using both ‘curated‘ and ‘all‘ sets. In Fig. 11, we show a control experiment in which we inpaint in-domain object class ‘car‘ into Cityscapes scenes, and we analyze models’ responses to synthetic cars, real cars, and OOD objects.
@@ 4.2 Assess OOD Detection
Figure 10: Qualitative results. Confidence maps are visualized for the four exemplified models on real data (first row) and synthetic inpainted data (second and third rows). Hotter colors correspond to higher OOD likelihood. Ideally, results should exhibit hot colors in OOD areas and cold colors everywhere else. We observe a strong correlation in model reactions to real and synthetic OOD regions, particularly for more recent and robust models. We note that the real data in SMIYC also exhibit distributional shifts against Cityscapes; which already causes confusion to weak models in the background.
Figure 11: In-domain vs OOD inpainted objects. Models’ responses to synthetic cars are close to real cars, and far from synthetic OOD objects
We then analyze the correlation between real and synthetic anomaly metrics.  Fig. 9 presents our primary findings and Fig. 10 illustrates some qualitative results. We observe a strong correlation with real scores when utilizing the ’curated’ set for computing synthetic scores; the curated PCCs ( ) are consistently around 0.8 across multiple metrics, irrespective of the two confidence measures. Although the correlations are somewhat weaker when using all uncurated synthetic data ( ), such results remain acceptable, particularly when no effort is dedicated to curation.
@@ 4.3 Improve OOD Detection
OOD metrics are computed on RoadAnomaly21 and reported in Tab. 2. The RbA models trained on our data significantly outperform the vanilla RbA model. We reach comparable performance to the RbA variants that leverage the external COCO dataset for augmentation. Notably, there are no clear differences between using ‘curated’ or ‘all’ sets. We conjecture that, unlike benchmarking, training for OOD detection does not demand a high degree of realism from synthetic data. This explains why the simple strategy of copy-pasting COCO objects [25] already proves effective. All results are consistent across the two addressed backbones.  Fig. 10 illustrates a few qualitative results.
Discussion. In our work, we take all available levers and study the extent to which synthetic data can be used for evaluation and specific training purposes. We acknowledge the fact that the models used in this work were trained on substantial amounts of data. Of note, we do not claim the efficacy of synthetic data in all aspects, particularly regarding the total amount of training data required. Our findings are limited to highlighting the significant potential of published generative models in the task of reliability assessment. Advances in efficient training of generative models may address concerns regarding data quantity but are beyond the scope of this work.
@@ 5 Takeaways
In this work, we explore the potential of synthetic data in reliability assessment for semantic segmentation networks. We introduce two automatic zero-shot pipelines to generate data in OOD domains and to inpaint OOD objects for virtual reliability assessment. Our promising results encourage further collective investigations into this research problem, paving the way for synthetic system validation, especially in safety-critical applications. We summarize here our findings:
@@
    Reliability Under Covariate Shifts: synthetic data can help assess the relative robustness of models in real-life covariate shifts, especially when shifts to the training condition are significant. Synthetic data can well complement real data in system validation, helping reduce the total operational cost. Pretrained models can be calibrated using synthetic data to better estimate prediction confidence in any shifted domains.
@@
Acknowledgements
This work is supported by ELSA - European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. We thank the authors of RELIS [14] for providing us their pretrained checkpoints.
In this document, we provide technical details for covariate shifts generation and OOD object inpainting (Appendix 0.A), additional calibration details and results (Appendix 0.B), analysis on direct real-vs.-synthetic data correlation (Appendix 0.C) and class-wise PCC scores (Appendix 0.D). Appendix 0.E discusses the limitations and Appendix 0.F showcases more qualitative examples. footnotetext: † Work done during an internship at valeo.ai
Appendix 0.A Technical Details
0.A.1 Covariate Shifts Training
@@
Similarly for SDXL, we train a ControlNet for 27500 steps and use a batch size of 8 with 4 gradient accumulation steps, which makes an effective batch size of 32. The learning rate is $10^{-5}$. The images are square cropped to get 1024 $\times$ 1024 images, convenient for SDXL, and there is no need to resize them.
0.A.2 Covariate Shifts Generation
To generate images with new styles with SD 1.5, we take a semantic mask from the validation set of Cityscapes, crop and resize it as explained in the previous part. We use nearest neighbor interpolation to keep good values for specific classes. We only use the part of the caption extracted with CLIP-interrogator that corresponds to a BLIP caption. To this new caption, we add [, in $<$domain$>$] depending on the domain we want to generalize to. Starting from pure noise, we use 25 DDIM steps with a guidance scale of 8. On a RTX 2080, one new image is generated in about 4 seconds. All other sampling hyperparameters are the per default settings on the official ControlNet repository.
For SDXL, we tune the conditioning strength and the prompt strength to get better images. To get both mask and prompt aligned images, we set the ControlNet conditioning strength to 0.65 and the prompt guidance to 10. The prompt tuning is the same as used for SD 1.5; we also use 25 denoising steps.
0.A.3 OOD Objects Generation
@@
The OOD object generation pipeline described in Sec. 4.1 is further illustrated in Fig. 12. It is composed of three steps. Given a text prompt containing an object, the inpainting step generates a zoomed-in version of the object with the appropriate close-range background given as context. The mask extraction step infers the anomaly mask from the zoomed-in generated image and the name of the object. Both are in-pasted back in the original complete image or mask. To reduce some composition artifacts, the composite image is refined with a noise/denoise step.
In details, we first randomly choose a box size for the new object, uniformly sampled between a quarter and half the minimum dimension of the original image. We also uniformly sample a location for the box in the bottom three quarters of the image. This box will contain the new object we wish to add, and we refer to it as inpainted region. In addition to the inpainted region, we create a larger box, 1.5$\times$ its height and width, with the inpainted region in its center. The contour outside of the inpainted region will serve as context for the inpainting process. We then crop and resize the inpainted region with its context to a resolution of 512 $\times$ 512. We fully noise the inpainted region, but leave the clean context. We then denoise the inpainted zone with the prompt “A photo of an [[[[object]]]], with a guidance scale of 15. The full patch is then resized and pasted on the original image, at its original position. Some artifact might be still present as shown in Fig. 13. To remedy this, we refine the inpainted zone by noising and denoising it with 0.65 strength, with the default guidance scale of 7.5. The effect of this refining step is shown in Fig. 13.
@@ 0.A.4 OOD Detection Training
For the OOD detection method in Sec. 4.2, we used the codebase of [25]. We adapt the code to be able to use our generated data with the binary masks extracted from Grounded-SAM, as explained in Sec. 0.A.3. As in the original paper [25], we fine-tune the mask prediction MLP and classification layer after the transformer decoder. To obtain all OOD detection results reported in the main paper (Tab. 2), we used the recommended hyperparameters, and train the models for 5000 iterations.
0.A.5 Segmentation Models